In a triangle ABC what would be the value of sinA+sinB-sinC?

Grade:11

2 Answers

sahil
6 years ago
in a triangle angle A+B+C=180(degree)so we need to find sin(a)+sin(b)+sin(c)=2sin(a+b/2).cos(a-b/2)+2sin(c/2)cos(c/2)=2sin(
sahil
6 years ago
in a triangle angle A+B+C=180(degree)so we need to find sin(a)+sin(b)-sin(c)=2sin(a/2+b/2).cos(a/2-b/2)-2sin(c/2)cos(c/2)=2cos(c/2).cos(a/2-b/2)-2sin(c/2)cos(c/2)=2cos(c/2)[cos(a/2-b/2)-cos(c/2)]=2cos(c/2)[cos(a/2-b/2)+cos(a/2+b/2)]=2cos(c/2)[cos(a/2).cos(b/2)]=4cos(a/2)cos(b/2).cos(c/2)